# Diffusion Models for Reinforcement Learning: A Survey

### 11 Dec 2023 v2

@@ 5 Applications of Diffusion Models
Diffusion models have recently been employed to address various problems in RL and several strongly related fields. We group these applications into four categories based on the task: offline RL, imitation learning, trajectory generation, and data augmentation. For each category, we provide a brief introduction to the task, followed by a detailed explanation of how existing works use diffusion models to improve performance on that task.
@@ 5.1 Offline RL
Offline RL Levine et al. (2020) aims to learn a policy from previously collected datasets without online interaction. Assuming there is a static dataset $\mathcal{D}$ collected by some (unknown) behavior policy $\pi_{\beta}$, offline RL requires the learning algorithm to derive a policy $\pi(a|s)$ that attains the most cumulative reward, which is defined in Eq. (10). The fundamental challenge in offline RL is the distributional shift: while the function approximator (e.g., policy, value function) might be trained under one distribution, it will be evaluated on a different distribution, leading to poor performance of the learned policy. High-dimensional and expressive function approximation generally exacerbates this issue:
$\pi^{*}:=\arg\max_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\Big{[}\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\Big{]}~{}.$ (10)
@@
Several methods use diffusion models to help tackle or avoid the above challenge. Janner et al. (2022) first propose to generate optimal trajectories through iteratively denoising with classifier-guided sampling. Subsequent works Wang et al. (2023); Chen et al. (2023a); He et al. (2023b); Ada et al. (2023); Brehmer et al. (2023); Hansen-Estruch et al. (2023); Venkatraman et al. (2023) represent the policy as a diffusion model to capture multi-modal distributions and enhance the expressiveness of the policy class, which is beneficial to relieve the approximation error between the cloned behavior policy and true behavior policy. Ajay et al. (2023) skip the risk of distribution shift by generating state sequences with conditional diffusion models followed by inverse dynamic functions to derive executable actions, which propose a novel approach to use classifier-free guidance with low-temperature sampling to denoise out return-maximizing trajectories. In order to improve the generation ability of diffusion models for RL, Lu et al. (2023a) propose a new guidance method named contrastive energy prediction and Hu et al. (2023) capture more temporal conditions. By incorporating control-theoretic invariance into the diffusion dynamics, SafeDiffuser Xiao et al. (2023) guarantees the safe generation of planning trajectories. HDMI Li et al. (2023a) leverages a hierarchical structure to tackle long-horizon decision-making problems, which uses a reward-conditional model to discover sub-goals and a goal-conditional model to generate actions. AlignDiff Dong et al. (2023) conditions on behavior attributes with classifier-free guidance to plan to match desired trajectories accurately. CPQL Chen et al. (2023d) leverages consistency models for fast training and sampling, while EDP Kang et al. (2023) obtains speed-up by using single-step model predictions as action approximations. Diffusion models can also be used to extract the reward function Nuti et al. (2023) or value function Mazoure et al. (2023), and are prominent to estimate gradients with score-matching to solve offline optimization problems Suh et al. (2023). Enabling RL agents to generalize to multi-task and multi-agent scenarios remains a challenge due to their inherent complexity. Recent research has made progress in using diffusion models to improve the performance of policies in multi-task and multi-agent offline RL.
@@
Diffusion model is verified to have the potential to address the challenge of multi-task generalization in RL. He et al. (2023a) first extend the conditional diffusion model to be capable of solving multitask decision-making problems and synthesizing useful data for downstream tasks. LCD Zhang et al. (2023a) leverages a hierarchical structure to achieve long-horizon multi-task control. Ni et al. (2023) and Liang et al. (2023) extend the idea of Diffuser Janner et al. (2022) into more specific sittings. MetaDiffuser Ni et al. (2023) demonstrates that incorporating the conditional diffusion model into the context of task inference outperforms previous meta-RL methods. AdaptDiffuser Liang et al. (2023) combines bootstrapping and diffusion-based generative modeling together to enable the model to adapt to unseen tasks.
@@
Employing diffusion models to multi-agent RL helps model discrepant behaviors among agents and reduces approximation error. MADiff Zhu et al. (2023) uses an attention-based diffusion model to model the complex coordination among behaviors of multiple agents, which is well-suited to learning complex multi-agent interactions. DOM2 Li et al. (2023c) incorporates the diffusion model into the policy classes to enhance learning and makes it possible to generalize to shifted environments well.
@@ 5.2 Imitation Learning
The goal of imitation learning (IL) is to reproduce behavior similar to experts in the environment by extracting knowledge from expert demonstrations. Recently, many works Hegde et al. (2023); Ng et al. (2023); Chen et al. (2023c); Kapelyukh et al. (2022) have demonstrated the efficacy of representing policies as diffusion models to capture multi-modal behavior. Pearce et al. (2023) apply diffusion models to imitate human behavior in sequential environments, in which diffusion models are compared with other generative models and viable approaches are developed to improve the quality of behavior sampled from diffusion models. Chi et al. (2023); Xian et al. (2023) generate the robot’s behavior via a conditional denoising diffusion process on robot action space. Experiment results show that Diffusion models are good at predicting closed-loop action sequences while guaranteeing temporal consistency Chi et al. (2023). Li et al. (2023b) improve the models in Chi et al. (2023) by incorporating an auxiliary reconstruction loss on intermediate representations of the reverse diffusion process. Beneficial from its powerful generation ability, leveraging diffusion models to acquire diverse skills to handle multiple manipulation tasks is promising Chen et al. (2023c); Mishra et al. (2023); Xu et al. (2023); Ha et al. (2023). Diffusion models are already applied to goal-conditioned RL: Reuss et al. (2023) use a decoupled score-based diffusion model to learn an expressive goal-conditional policy. In contrast, Sridhar et al. (2023) build a unified diffusion policy to solve both goal-directed navigation and goal-agnostic exploration problems.
@@ 5.3 Trajectory Generation
Trajectory generation aims to derive a dynamically feasible path that satisfies a set of constraints. In particular, we focus on generating human pose and robot interaction sequences, which are more related to the decision-making scenario. Many works Zhang et al. (2022); Jiang et al. (2023); Tevet et al. (2022); Zhang et al. (2023b); Chen et al. (2022); Dabral et al. (2022) have remarked that the conditional diffusion models perform better than traditional methods which use GAN or Transformer. Employing a denoising-diffusion-based framework, they achieve diverse and fine-grained motion generation with various conditioning contexts Chen et al. (2023b); Carvalho et al. (2023). Recent works Du et al. (2023b); Ko et al. (2023); Du et al. (2023a) harness diffusion models to synthesize a set of future frames depicting its planned actions in the future, after which control actions are extracted from the generated video. This approach makes it possible to train policies solely based on RGB videos and deploy learned policies to various robotic tasks Black et al. (2023); Gao et al. (2023). UniSim Yang et al. (2023) uses diffusion models to build a universal simulator of real-world interaction by learning through combined diverse datasets. It can be utilized to train both high-level vision-language planners and low-level RL policies, demonstrating powerful emulation ability.
@@ 5.4 Data Augmentation
Diffusion models have already been verified to be useful for data augmentation in the RL domain. Since diffusion models perform well in learning over multimodal or even noisy distributions, they can model original data distribution precisely. What is more, they are capable of generating diverse data points to expand original distribution while maintaining dynamic accuracy. Recent works Yu et al. (2023); Chen et al. (2023e) consider augmenting the observations of robotic control using a text-guided diffusion model while maintaining the same action. The recently proposed SynthER Lu et al. (2023b) and MTDiff-s He et al. (2023a) generate complete transitions of trained tasks via a diffusion model. It proves that such augmentation brings about significant policy improvement for both online and offline RL.
6 Limitations and Remarks
In this section, we list three limitations when applying diffusion models in RL, and include remarks on comparing diffusion-based generative modeling in RL to transformer-based autoregressive approaches.
Application in online RL.
Though diffusion models have contributed to offline RL from various perspectives, the dynamic and evolving nature of online RL introduces a significant challenge, as the data distribution sampled by the current policy can change over time. The primary concern is that effectively adapting diffusion models to changing data distributions requires a substantial amount of new data Lu et al. (2023b). These models are typically trained on fixed datasets, and incorporating enough data to ensure they can generalize and remain effective in dynamic real-world scenarios is a resource-intensive task. Balancing the need for adaptability with the requirement for extensive data is the primary consideration when applying diffusion models in online RL. It is promising to solve this dilemma with more lightweight diffusion models which can keep consistency as the data distribution changes during online interactions.
Iterative sampling cost.
The unique stepwise denoising mechanism of diffusion models makes the sampling procedure need to infer the trained model multiple times, significantly increasing the sampling cost. Although we can perform sampling acceleration techniques such as DDIM or DPM-Solver, the expensive inference cost limits the model to conduct high-frequency output in online interactions with virtual environments and real-world continuous control scenarios. The problem becomes more severe in those methods that generate a long trajectory but only take the first state or action for execution, such as Diffuser Janner et al. (2022) or Decision Diffuser Ajay et al. (2023). Chen et al. (2023d) incorporate the recently proposed Consistency Model Song et al. (2023) to enable sampling with one or two diffusion steps, and the performances are comparable to that of DDPM or DDIM with 50 steps.
Variance in stochastic sampling.
In traditional RL algorithms, a continuous control policy is represented by a state-conditioned Gaussian Haarnoja et al. (2018); Schulman et al. (2017). We can take its mean as action when deterministic execution is required. However, such deterministic policy is not possible when using diffusion models as the policy class. The randomness of diffusion sampling comes from both initial noise and per-step stochastic denoising. Although per-step stochasticity can be avoided if sampling with ODE-based methods like DDIM, the randomness in initial noisy samples remains inevitable. The high-variance policies can have a negative impact in environments with high accuracy or safety requirements. Existing works in RL rarely discuss this limitation, and sampling methods with reduced variances are expected.
Comparison to transformer-based methods.
The idea of using diffusion models commonly abstracts offline RL as a conditional generative modeling problem, differing from traditional RL approaches, which need value function approximations or policy gradient calculation. This supervised learning paradigm is similar to Decision Transformer (DT) Chen et al. (2021), a sequential modeling framework for RL tasks with transformer architecture. The essence of both approaches is to take advantage of the high expressiveness model, where diffusion models enjoy a strong distribution fitting ability to produce multi-modal, diverse, and accurate outputs, and transformers are adept in long-horizon sequence modeling and time correlation understanding. Qualitatively, this difference in expertise makes the diffusion-based approach more suitable for learning complex multi-modal tasks, and Transformer-based approaches Meng et al. (2021); Wen et al. (2022) are more preferred in correlated sequence modeling in time or agent (in multi-agent tasks) dimension.
@@ 7 Summary and Future Prospects
This survey offers a comprehensive overview of contemporary research endeavors concerning the application of diffusion models in the realm of RL. According to the roles played by diffusion models, we categorize existing methodologies into using diffusion models as planners, policies, data synthesizers, and less popular roles such as value functions, latent representation models, etc. By comparing each class of methods to traditional solutions, we can see how the diffusion model addresses some of the longstanding challenges in RL, i.e., restricted expressiveness, data scarcity, compounding error, and multitask generalization. Notwithstanding these merits, it is imperative to acknowledge the existence of non-negligible limitations in using diffusion models in RL due to some inherent properties in the training and sampling of diffusion models. It is worth emphasizing that the incorporation of diffusion models into RL remains an emerging field, and there are many research topics worth exploring. Here, we outline three prospective research directions, namely, retrieval-enhanced generation, integrating safety constraints, and composing different skills.
@@
Integrating safety constraints.
Utilizing RL models for practical applications often necessitates compliance with various safety constraints García and Fernández (2015); Gu et al. (2022). Several safe RL methods transform a constrained RL problem to its unconstrained equivalent Achiam et al. (2017); Tessler et al. (2018); Sootla et al. (2022), which is then solved by generic RL algorithms. Policies acquired through these methods remain tailored to the specific constraint threshold utilized during training. Recent research Liu et al. (2023); Zhang et al. (2023c) has extended the applicability of decision transformers to the context of safety-constrained settings, thereby enabling a single model to adapt to diverse thresholds by adjusting the input cost-to-go. Similarly, diffusion models exhibit substantial potential for deployment in the domain of safe RL. Ajay et al. (2023) demonstrate that a diffusion-based planner can combine different movement skills to produce new behaviors. In addition, classifier-guided sampling can add new conditions to generated samples simply by learning additional classifiers, while the parameters of the diffusion model itself remain unchanged Dhariwal and Nichol (2021). This feature makes the diffusion model promising to be effective in scenarios where new safety constraints can be included after model training.
@@
Most current works deploy the generation ability of diffusion models on the raw state and action spaces. From the perspective of skill-based RL Shi et al. (2022); Nam et al. (2022), it is promising to break down complex tasks into smaller, more manageable sub-skills. Diffusion models excel in modeling multi-modal distributions, and since multiple sub-skills can be viewed as distinct modes within the distribution of possible behaviors, they offer a natural fit for this task. Combining with classifier guidance or classifier-free guidance, diffusion models are possible to generate proper skills to complete the facing task. Experiments in offline RL also suggest that diffusion models can share knowledge across skills and combine them up Ajay et al. (2023); He et al. (2023a), thus having the potential for zero-shot adaptation or continuous RL by composing different skills.

### 23 Feb 2024 v4

@@ 5 Applications of Diffusion Models
In this section, we conduct a complete review of existing works. We divide them into five groups based on the tasks they are applied: offline RL, online RL, imitation learning, trajectory generation, and data augmentation. For each group, we provide a detailed explanation of how each method uses diffusion models to handle the task.
@@ 5.1 Offline RL
Offline RL aims to learn a policy from previously collected datasets without online interaction. Assuming there is a static dataset $\mathcal{D}$ collected by some (unknown) behavior policy $\pi_{\beta}$, offline RL requires the learning algorithm to derive a policy $\pi(a|s)$ that attains the most cumulative reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\Big{[}\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\Big{]}~{}.$
The fundamental challenge in offline RL is the distributional shift. This refers to the discrepancy between the dataset distribution used to train the function approximators (e.g., policies and value functions) and the distribution on which the policy is evaluated. This mismatch often results in subpar online performance. High-dimensional and expressive function approximation generally exacerbates this issue.
@@
Several methods use diffusion models to help tackle or avoid the above challenges. Janner et al. (2022) first propose to generate optimal trajectories through iterative denoising with classifier-guided sampling. Subsequent works Wang et al. (2023); Chen et al. (2023a); He et al. (2023b); Ada et al. (2023); Brehmer et al. (2023); Hansen-Estruch et al. (2023) represent the policy as a diffusion model to capture multi-modal distributions and enhance the expressiveness of the policy class, which is beneficial to relieve the approximation error between the cloned behavior policy and true behavior policy. Ajay et al. (2023) alleviate the distribution shift problem by generating state sequences with conditional diffusion models followed by inverse dynamic functions to derive executable actions, which propose a novel approach to use classifier-free guidance with low-temperature sampling to denoise out return-maximizing trajectories. LatentDiffuser Li (2023) performs diffusion planning over a learned latent space with separate decoders to recover raw trajectories. Benefiting from a more compact planning space, it achieves superior performances on long-horizon and high-dimensional tasks. In order to improve the generation ability of diffusion models for RL, Lu et al. (2023a) propose a new guidance method named contrastive energy prediction and Hu et al. (2023) capture more temporal conditions. By incorporating control-theoretic invariance into the diffusion dynamics, SafeDiffuser Xiao et al. (2023) guarantees the safe generation of planning trajectories. HDMI Li et al. (2023a) adopts a hierarchical structure to tackle long-horizon decision-making problems, which uses a reward-conditional model to discover sub-goals and a goal-conditional model to generate actions. Dong et al. (2023) condition the diffusion planner on diverse behavior attributes and learn from human preferences to generate trajectories that can match user-customized behaviors. CPQL Chen et al. (2023d) leverages consistency models as the policy class for fast training and sampling, while EDP Kang et al. (2023) achieves speed-up during training by using single-step model predictions as action approximations. Diffusion models are also used as the value function Mazoure et al. (2023) and representation model Venkatraman et al. (2023) to facilitate training of normal RL policies. Recent research has made progress in using diffusion models to improve the performance of policies in multitask and multi-agent offline RL.
@@
Diffusion models are verified to have the potential to address the challenge of multi-task generalization in offline RL. He et al. (2023a) first extend the conditional diffusion model to be capable of solving multitask decision-making problems and synthesizing useful data for downstream tasks. LCD Zhang et al. (2023a) leverages a hierarchical structure to achieve long-horizon multi-task control. MetaDiffuser Ni et al. (2023) demonstrates that incorporating the conditional diffusion model into the context of task inference outperforms previous meta-RL methods. AdaptDiffuser Liang et al. (2023a) combines bootstrapping and diffusion-based generative modeling together to enable the model to adapt to unseen tasks.
@@
Using diffusion models in multi-agent RL helps model discrepant behaviors among agents and reduces approximation error. MADiff Zhu et al. (2023) uses an attention-based diffusion model to model the complex coordination among behaviors of multiple agents, which is well-suited to learning complex multi-agent interactions. DOM2 Li et al. (2023c) incorporates the diffusion model into the policy classes to enhance learning and makes it possible to generalize to shifted environments well.
5.2 Online RL
Recently, there have been some works showing that diffusion models can also boost online RL training. Value estimations in online RL are noisier and change with the current policy, which poses additional challenges on training a multistep diffusion model. DIPO Yang et al. (2023a) proposes an action relabeling strategy to perform policy improvement at the data level, bypassing the potentially unstable value-guided training. Actions in the online rollout dataset are updated with gradient ascent, and the diffusion training objective is just supervised learning on the relabeled dataset. Chen et al. (2023d) conduct experiments to verify that consistency models with one-step sampling can naturally serve as online RL policies and achieve a balance between exploration and exploitation. Instead of using diffusion models as policies, Rigter et al. (2023) build a diffusion dynamic model to generate synthetic trajectories that are consistent with online RL policies. Applications of diffusion models in online RL are less explored compared to offline settings and merit further investigation.
@@ 5.3 Imitation Learning
The goal of imitation learning (IL) is to reproduce behavior similar to experts in the environment by extracting knowledge from expert demonstrations. Recently, many works Hegde et al. (2023); Ng et al. (2023); Chen et al. (2023c); Kapelyukh et al. (2022) have demonstrated the efficacy of representing policies as diffusion models to capture multi-modal behavior. Pearce et al. (2023) apply diffusion models to imitate human behavior in sequential environments, in which diffusion models are compared with other generative models and viable approaches are developed to improve the quality of behavior sampled from diffusion models. Chi et al. (2023); Xian et al. (2023) generate the robot’s behavior via a conditional denoising diffusion process on robot action space. Experiment results show that Diffusion models are good at predicting closed-loop action sequences while guaranteeing temporal consistency Chi et al. (2023). Li et al. (2023b) improve the models in Chi et al. (2023) by incorporating an auxiliary reconstruction loss on intermediate representations of the reverse diffusion process. Beneficial from its powerful generation ability, leveraging diffusion models to acquire diverse skills to handle multiple manipulation tasks is promising Chen et al. (2023c); Mishra et al. (2023); Xu et al. (2023); Ha et al. (2023). Diffusion models are already applied to goal-conditioned IL: Reuss et al. (2023) use a decoupled score-based diffusion model to learn an expressive goal-conditional policy. In contrast, Sridhar et al. (2023) build a unified diffusion policy to solve both goal-directed navigation and goal-agnostic exploration problems. Liang et al. (2023b) adopt a hierarchical structure where the high-level skills are determined by the current visual observation and language instructions. Therefore the low-level skill-conditioned planner can satisfy the user-specified multitask instructions.
@@ 5.4 Trajectory Generation
Trajectory generation aims to produce a dynamically feasible path that satisfies a set of constraints. We focus on using diffusion models to generate human pose and robot interaction sequences, which are more related to the decision-making scenario. Many works Zhang et al. (2022); Jiang et al. (2023); Tevet et al. (2022); Zhang et al. (2023b); Chen et al. (2022); Dabral et al. (2022) have remarked that the conditional diffusion models perform better than traditional methods which use GAN or Transformer. Employing a denoising-diffusion-based framework, they achieve diverse and fine-grained motion generation with various conditioning contexts Chen et al. (2023b); Carvalho et al. (2023). Recent works Du et al. (2023b); Ko et al. (2023); Du et al. (2023a) harness diffusion models to synthesize a set of future frames depicting its planned actions in the future, after which control actions are extracted from the generated video. This approach makes it possible to train policies solely on RGB videos and deploy learned policies to various robotic tasks with varying dynamics Black et al. (2023); Gao et al. (2023). UniSim Yang et al. (2023b) uses diffusion models to build a universal simulator of real-world interaction by learning through combined diverse datasets. It can be used to train both high-level vision-language planners and low-level RL policies, demonstrating powerful emulation ability.
@@ 5.5 Data Augmentation
Since diffusion models perform well in learning over multimodal or even noisy distributions, they can model original data distribution precisely. What is more, they are capable of generating diverse data points to expand original distribution while maintaining dynamic accuracy. Recent works Yu et al. (2023); Chen et al. (2023e) consider augmenting the observations of robotic control using a text-guided diffusion model while maintaining the same action. The recently proposed SynthER Lu et al. (2023b) and MTDiff-s He et al. (2023a) generate complete transitions of trained tasks via a diffusion model. Lu et al. (2023b) directly train the diffusion model from the offline dataset or the online replay buffer and then generate samples for policy improvement. Analysis shows that both diversity and accuracy of data generated by diffusion models are higher than those generated by prior data augmentation methods. He et al. (2023a) deploy diffusion synthesizer on multi-task offline datasets and achieve better performance than that on single-task datasets. They claim that fitting on multiple tasks may enable implicit knowledge sharing across tasks, which also benefits from the multi-modal property of diffusion models. These works demonstrate that diffusion augmentation can bring significant improvement in policy learning for IL, online RL, and offline RL.
@@ 6 Summary and Future Prospects
This survey offers a comprehensive overview of contemporary research endeavors concerning the application of diffusion models in the realm of RL. According to the roles played by diffusion models, we categorize existing methods into using diffusion models as planners, policies, data synthesizers, and less popular roles such as value functions, representation models, etc. By comparing each class of methods to traditional solutions, we can see how the diffusion model addresses some of the longstanding challenges in RL, i.e., restricted expressiveness, data scarcity, compounding error, and multitask generalization. It is worth emphasizing that the incorporation of diffusion models into RL remains an emerging field, and there are many research topics worth exploring. Here, we outline four prospective research directions, namely, generative simulation, integrating safety constraints, retrieval-augmented generation, and composing different skills.
Generative simulation.
As shown in Fig. 1, existing works use diffusion models to overcome certain limitations of previous solutions in both the agent and buffer parts. However, there has been a scarcity of research focused on using diffusion modeling to improve the environment. Gen2Sim Katara et al. (2023) uses text-to-image diffusion models to generate diverse objects in simulation environments, where RL policies are trained to learn robot manipulation skills. Besides generating objects in the scene, diffusion models have the potential for broader applications in generative simulation, such as the generation of various possible dynamics functions, reward functions, or opponents in multi-agent learning.
Integrating safety constraints.
Making decisions in real tasks often necessitates compliance with various safety constraints. Several safe RL methods transform a constrained RL problem to its unconstrained equivalent Achiam et al. (2017), which is then solved by generic RL algorithms. Policies acquired through these methods remain tailored to the specific constraint threshold specified during training. A recent research Liu et al. (2023) has extended the applicability of decision transformers to the context of safety-constrained settings, thereby enabling a single model to adapt to diverse thresholds by adjusting the input cost-to-go. Similarly, diffusion models have the potential to be deployed in safe RL by viewing safety constraints as sampling conditions. Ajay et al. (2023) demonstrate that a diffusion-based planner can combine different movement skills to produce new behaviors. Also, classifier-guided sampling can include new conditions by learning additional classifiers, while the parameters of the diffusion model remain unchanged Dhariwal and Nichol (2021). This makes the diffusion model promising for scenarios with new safety requirements after model training.
@@
From the perspective of skill-based RL, it is promising to break down complex tasks into smaller, more manageable sub-skills. Diffusion models excel in modeling multi-modal distributions, and since multiple sub-skills can be viewed as distinct modes within the distribution of possible behaviors, they offer a natural fit for this task. Combining with classifier guidance or classifier-free guidance, diffusion models are possible to generate proper skills to complete the facing task. Experiments in offline RL also suggest that diffusion models can share knowledge across skills and combine them up Ajay et al. (2023), thus having the potential for zero-shot adaptation or continuous RL by composing different skills.